OpenStudy (anonymous):

The state income tax in Connecticut is modeled by the function below, where x is income in dollars. Suppose that Connecticut decided to increase its tax rates by 20% and add a filing fee of $100. Write a function for the new state income tax. f(x) = {0.02 x if 010,000

OpenStudy (anonymous):

f(x)= (0.02+0.2)x+100

OpenStudy (anonymous):

what about the rest of the piecewise ?

OpenStudy (anonymous):

same thing i think

OpenStudy (zepp):

That sounds right

OpenStudy (zepp):

Actually

OpenStudy (zepp):



OpenStudy (zepp):

First one is right f(x)= (0.02+0.2)x+100 and second one should following the same pattern

OpenStudy (zepp):

y=0.05x would become y = (0.05+0.2)x+100

OpenStudy (zepp):

Put them together f(x) = (0.02+0.2)x+100 for 0<x<= 10,000 (0.05+0.2)x+100 for x>10,000

OpenStudy (anonymous):

oh for thre <=10,000 its suppose to be less than or equal to

OpenStudy (zepp):

I know

OpenStudy (anonymous):

these are two of the multiple choice answers I was given there are two more

OpenStudy (anonymous):

heres the other 2, this is why Im confused on your guys answers

OpenStudy (anonymous):

I dont get how you guys got your answers